OpenStudy (anonymous):

What expression is in front of the radical in your simplified answer? 3a sqrt.b^3 x 4 sqrt a 12ab?

OpenStudy (kropot72):

The simplified form is as follows: \[3a \sqrt{b ^{3}}\times4\sqrt{a}=12a \sqrt{ab ^{3}}\]

OpenStudy (anonymous):

12 a logical thank u

OpenStudy (kropot72):

You're most welcome:)